Find the potential at a distance of 0.07m from a -6nC charge.





1
Expert's answer
2022-03-09T12:30:52-0500

Potential

"V=\frac{kQ}{r}"

"V=-\frac{9\times10^9\times6\times10^{-9}}{0.07}=-771.42C"